OpenStudy (vera_ewing):

Choose the graph that matches the following system of equations: 8x + 5y = –5 7x – 3y = 3

OpenStudy (vera_ewing):

@amistre64 Is it C?

OpenStudy (texaschic101):

8x + 5y = -5 5y = -8x -5 y = -8/5x - 1 slope is -8/5.....y intercept is (0,-1) x intercept can be found by subbing in 0 for y 0 = -8/5x - 1 8/5x = -1 x = -1/(8/5) x = -1 * 5/8 x = -5/8......so your x intercept is (-5/8,0) 7x - 3y = 3 -3y = -7x + 3 y = 7/3x - 1 slope is 7/3.....y intercept is (0,-1) x int : 0 = 7/3x - 1 -7/3x = -1 x = -1/(-7/3) x = -1 * -3/7 x = 3/7......so your x intercept is (3/7,0) yes...it is C

OpenStudy (vera_ewing):

thanks so much :)

OpenStudy (texaschic101):

sure thing :)